OpenStudy (anonymous):

First factor the trinomial as a binomial square and then use the square root method to solve. Round the solution to the nearest tenth. x2 - 14x + 49 = 3 A. {5.3, 8.7} B. {4.8, 9.2} C. {3.5, 7.5} D. {6.2, 8.3} First factor the trinomial as a binomial square and then use the square root method to solve. Round the solution to the nearest tenth. x2 - 14x + 49 = 3 A. {5.3, 8.7} B. {4.8, 9.2} C. {3.5, 7.5} D. {6.2, 8.3} @Mathematics

OpenStudy (anonymous):

(x-7)^2 = +isqrt3 x = sqrt3 + 7 = 8.7 or x = -sqrt3 + 7 = 5.3 A

OpenStudy (anonymous):

Thank you